OpenStudy (anonymous):

Show that i^i is real, where i = sqrt(-1).

OpenStudy (anonymous):

Let i^i = e^x -> i ln i = x i ln (sqrt(-1)) = x i/2 ln -1 = x but since ln -1 = ipi x = -pi/2